A Tale of Two Applications on the NII

Geoffrey C. Fox Syracuse University Northeast Parallel Architectures Center 111 College Place Syracuse, New York 13244-4100 gcf@npac.syr.edu, http://www.npac.syr.edu

Abstract

We describe the expected capability of the NII (as an evolution of the Internet) interns of five broad service areas—collaboration, multimedia information dissemination, commerce, metacomputing, and Webtop productivity. We illustrate the demands on these services and the technology implications by examination of two application areas— manufacture of complex systems, such as aircraft and crisis management, command and control.